OpenStudy (anonymous):

Solve the equation. Give an approximate solution to four decimal places :: e(x + 4) = 6 A) 2.3026 B) -2.1282 C) -2.2082 D) -1.6974

OpenStudy (lilg132):

e(x+4) = 6 are you sure you wrote that correct?

OpenStudy (anonymous):

yeah. i just double checked and that's what it says....

OpenStudy (lilg132):

expanding it would give you ex + e4 = 6

OpenStudy (anonymous):

do not expand

OpenStudy (anonymous):

divide both sides by e, then minus 4

OpenStudy (lilg132):

so you have x + 4 = 6/e x = 2/e ?

OpenStudy (anonymous):

but I think the problem should be e^(x+4)=6

OpenStudy (anonymous):

x+4 should be the power of e

OpenStudy (anonymous):

okay. it doesnt look like that in the document i was sent but i think you're right.

OpenStudy (hoblos):

C) -2.2082

OpenStudy (anonymous):

hoblos, how did you get that?

OpenStudy (hoblos):

e(x+4) = 6 => x+4 = ln6 => x = ln6 - 4 = -2.2082

OpenStudy (hoblos):

btw e(x+4) is e^(x+4)

OpenStudy (anonymous):

THANKS!

OpenStudy (hoblos):

welcome :D

OpenStudy (anonymous):

\[e^{x+4}=6\] \[x+4=\ln(6)\] \[x=\ln(6)-4\] then a calculator